It’s been suggested that we use “3” as an estimate for the value of π. Using both “3” and “3.14159”, calculate the volume and surface area of the earth. Then determine the percent difference between the methods.

8. Its been suggested that we use “3” as an estimate for the value of t. Using both “3” and “3.14159, calculate the volume
Answer #1


Ra bauxioom 3 R3 = 4*3*16.4410033 | V3 = 1.04858x1021 m3 l V3.14159 = 4* 3.1459x (604x1063 13-14159 = 1.09804 x 10.2163 % dil